OpenStudy (11study11):

Help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Algebra multiple choice.

OpenStudy (11study11):

@jim_thompson5910

OpenStudy (11study11):

A company dyes two sizes of rugs. A small rug requires 2 hours for dyeing and a medium-size rug requires 3 hours for dyeing. The dyers need to make at least 15 rugs, and they must do it in less than 60 hours. Let x equal small rugs and y equal medium rugs. Which of the following inequalities can be paired with x + y ≥ 15 to create a system that represents this situation? 2x + 3y < 60 3x + 2y < 60 2x + 3y > 60 3x + 2y > 60

OpenStudy (11study11):

A correct

OpenStudy (11study11):

@mathmate

OpenStudy (11study11):

@zepdrix

zepdrix (zepdrix):

2 hours per small rug: 2x 3 hours per medium rug: 3y The sum of these time values should be less than 60 hours. \(\large\rm 2x+3y<60\) Yes, good job!

OpenStudy (11study11):

Thanks! Can you help with some more?

OpenStudy (11study11):

Determine the solution to f(x) = g(x) using the following system of equations: f(x) = 5.5x − 13 g(x) = −5x + 18.5 x = 1 x = 2 x = 3 x = 4

zepdrix (zepdrix):

\[\large\rm f(x)\qquad~~=\qquad g(x)\]\[\large\rm 5.5x-13=-5x+18.5\]Solve for x. Or, maybe easier, plug in the options and see which one works.

zepdrix (zepdrix):

\[\large\rm 5.5\color{orangered}{x}-13=-5\color{orangered}{x}+18.5\]If we plug in x=1,\[\large\rm 5.5\color{orangered}{(1)}-13=-5\color{orangered}{(1)}+18.5\]Does the equation hold true?

OpenStudy (11study11):

no

OpenStudy (11study11):

3

zepdrix (zepdrix):

yay good job

OpenStudy (11study11):

The graph below shows the solution to a system of inequalities: Solid line joining ordered pairs 0, 3 and 7, 0. Shade the portion of the graph above the line in the first and fourth quadrant Which of the following inequalities is modeled by the graph? 2x + 5y ≥ 14; x ≥ 0 2x + 5y ≤ 14; x ≥ 0 2x − 5y ≤ 14; x ≥ 0 −2x − 5y ≤ 14; x ≥ 0 @zepdrix